# Symmetry-Preserving Finite-Difference Schemes and Auto-Bäcklund Transformations for the Schwarz Equation

### v2

@@ 1 Inroduction
$y^{\prime\prime}+2\,\frac{y^{\prime}+C{y^{\prime}}^{3/2}+{y^{\prime}}^{2}}{x-y}=0,$ (1.4)
where $C$ is an arbitrary constant. Notice that this constant cannot be eliminated by any equivalence transformation, i.e., without altering the group structure of the equation. This second-order ODE belongs to the Lie classification list [10, 11]. In his classification, Lie identified all second-order ODEs that admit nontrivial symmetries. Remarkably, most second-order ODEs whose solutions and integrals appear in standard handbooks are equivalent, up to point transformations, to equations from the Lie list.
@@
$y=\frac{1}{A(B-Ax)}+\frac{B-C}{A},$ (1.6)
where $A$ and $B$ are constants of integration. Equation (1.4) also has the singular solution
$y=ax+b,$ (1.7)
where $a$ and $b$ are constants, and $a$ must satisfy the condition
$a(a+C\sqrt{a}+1)=0.$ (1.8)
Notice that (1.6) is a particular case of the solution (1.3). Indeed, a connection between (1.1) and (1.4) can be established. [^1] For the case $C=0$, the connection between the solutions of these ODEs was noted long ago; see, e.g., [12, p. 15], which references the Mathematical Tripos, Part I Examination, 1911. By solving (1.4) for the constant $C$, one obtains
$\frac{(y-x)y^{\prime\prime}-2y^{\prime}\left(1+y^{\prime}\right)}{2{y^{\prime}}^{3/2}}=C.$ (1.9)
@@
$\theta\left(\frac{\sqrt{u_{\bar{x}}}}{\sqrt{(x-u_{-})(x_{-}-u)}}-\frac{x_{+}-u}{x_{-}-u}\frac{\sqrt{u_{x}}}{\sqrt{(x-u_{+})(x_{+}-u)}}\right)=\frac{C\sqrt{1+\varepsilon}\,(x-x_{-})u_{\bar{x}}}{(x-u_{-})(x_{-}-u)},$ (2.6)
$\frac{(x_{+}-x)^{2}u_{x}}{(x-u_{+})(x_{+}-u)}=\frac{(x-x_{-})^{2}u_{\bar{x}}}{(x-u_{-})(x_{-}-u)}=\varepsilon,$ (2.7)
where the small parameter $\varepsilon$ characterizes the density of the finite-difference mesh, and $\theta$ ($\theta\to 1$) is a parameter. In case
$\theta=\frac{\sqrt{1+\varepsilon}}{2}\left(|C|\sqrt{\varepsilon}+\sqrt{\varepsilon C^{2}+4}\right).$ (2.8)
the scheme becomes exact, [^2] To simplify the calculations, in the next section we assume that $\theta=\textrm{sgn}\,{C}$. As verified through previous calculations carried out by the authors, this assumption does not qualitatively affect the subsequent results, while it substantially simplifies the equations. i.e., its solutions at the grid nodes coincide with those of the corresponding ODE. As scheme (2.6), (2.7) is invariant, it admits the same algebra (1.5) as that of (1.4).
@@
$\begin{array}[]{c}\displaystyle J_{1}=\frac{u_{x}+1}{\sqrt{u_{x}(x-u)(x_{+}-u_{+})}}+\frac{C}{x_{+}-u}=A,\\ \displaystyle J_{2}=\frac{x_{+}u_{x}+u}{\sqrt{u_{x}(x-u)(x_{+}-u_{+})}}+\frac{Cx_{+}}{x_{+}-u}=B,\\ \displaystyle J_{3}=\frac{(x_{+}-x)(u_{+}-u)}{(x-u_{+})(x_{+}-u)}=\varepsilon.\end{array}$ (2.9)
@@
where $\rho$ is constant.
@@
$x_{n+1}=\frac{1}{2a(\varepsilon+1)}\left\{\left((a-1)\sqrt{\varepsilon(a^{2}\varepsilon+2(\varepsilon+2)a+\varepsilon)}+\varepsilon+(\varepsilon a+2)a\right)x_{n}\right.\\ \left.+\,b\left(\sqrt{\varepsilon(a^{2}\varepsilon+2(\varepsilon+2)a+\varepsilon)}+\varepsilon(a-1)\right)\right\}.$
This is a linear equation solvable by standard methods. Substituting this into (2.6) yields
$a^{2}C\sqrt{2(\varepsilon+1)}\sqrt{(a-1)\sqrt{\varepsilon(\varepsilon a^{2}+2(\varepsilon+2)a+\varepsilon)}+\varepsilon+(\varepsilon a+2)a}\\ -\theta a(a+1)\left(\sqrt{\varepsilon(a^{2}\varepsilon+2(\varepsilon+2)a+\varepsilon)}-\varepsilon-(\varepsilon+2)a\right)=0.$ (2.15)
Taking into account (2.8), in the continuous limit $(\varepsilon\to 0)$, one gets
$a\,(a+C\sqrt{a}+1)+O(\sqrt{\varepsilon})=0,$
which corresponds to the condition (1.8). This can also be interpreted as a condition on the choice of $\varepsilon$: for the scheme to have the singular solution (2.14), the mesh density parameter $\varepsilon$ must satisfy (2.15).
@@ 3 Establishing a Finite-Difference Connection Through Bäcklund-type Transformations (Case $K=C^{2}=4$)
$\frac{(x-u_{-})(x_{-}-u)}{(u-u_{-})\sqrt{1+\varepsilon}}\left(\frac{\sqrt{u_{\bar{x}}}}{\sqrt{(x-u_{-})(x_{-}-u)}}-\frac{x_{+}-u}{x_{-}-u}\frac{\sqrt{u_{x}}}{\sqrt{(x-u_{+})(x_{+}-u)}}\right)=C.$ (3.1)
@@
$\frac{(u-x_{-})(u_{-}-x)}{(x-x_{-})\sqrt{1+\varepsilon}}\left(\frac{\sqrt{x_{\bar{u}}}}{\sqrt{(u-x_{-})(u_{-}-x)}}-\frac{u_{+}-x}{u_{-}-x}\frac{\sqrt{x_{u}}}{\sqrt{(u-x_{+})(u_{+}-x)}}\right)=\widetilde{C},$ (3.2)
@@
$x_{u}=\frac{x_{+}-x}{u_{+}-u},\qquad x_{\bar{u}}=\frac{x-x_{-}}{u-u_{-}},\qquad\widetilde{C}=\frac{C}{1+C\sqrt{\varepsilon(1+\varepsilon)}}.$

### v3

@@ 1 Inroduction
$y^{\prime\prime}+2\,\frac{y^{\prime}+C_{0}{y^{\prime}}^{3/2}+{y^{\prime}}^{2}}{x-y}=0,$ (1.4)
where $C_{0}$ is an arbitrary constant. Notice that this constant cannot be eliminated by any equivalence transformation, i.e., without altering the group structure of the equation. This second-order ODE belongs to the Lie classification list [10, 11]. In his classification, Lie identified all second-order ODEs that admit nontrivial symmetries. Remarkably, most second-order ODEs whose solutions and integrals appear in standard handbooks are equivalent, up to point transformations, to equations from the Lie list.
@@
$y=\frac{1}{A_{0}(B_{0}-A_{0}x)}+\frac{B_{0}-C_{0}}{A_{0}},$ (1.6)
where $A_{0}$ and $B_{0}$ are constants of integration. Equation (1.4) also has the singular solution
$y=a_{0}x+b_{0},$ (1.7)
where $a_{0}$ and $b_{0}$ are constants, and $a_{0}$ must satisfy the condition
$a_{0}(a_{0}+C_{0}\sqrt{a}_{0}+1)=0.$
Notice that (1.6) is a particular case of the solution (1.3). Indeed, a connection between (1.1) and (1.4) can be established. [^1] For the case $C_{0}=0$, the connection between the solutions of these ODEs was noted long ago; see, e.g., [12, p. 15], which references the Mathematical Tripos, Part I Examination, 1911. By solving (1.4) for the constant $C_{0}$, one obtains
$\frac{(y-x)y^{\prime\prime}-2y^{\prime}\left(1+y^{\prime}\right)}{2{y^{\prime}}^{3/2}}=C_{0}.$ (1.8)
@@
$\theta\left(\frac{\sqrt{u_{\bar{x}}}}{\sqrt{(x-u_{-})(x_{-}-u)}}-\frac{x_{+}-u}{x_{-}-u}\frac{\sqrt{u_{x}}}{\sqrt{(x-u_{+})(x_{+}-u)}}\right)+\frac{C\sqrt{1+\varepsilon}\,(x-x_{-})u_{\bar{x}}}{(x-u_{-})(x_{-}-u)}=0,$ (2.6)
$\frac{(x_{+}-x)(u_{+}-u)}{(x-u_{+})(x_{+}-u)}=\frac{(x-x_{-})(u-u_{-})}{(x-u_{-})(x_{-}-u)}=\varepsilon,$ (2.7)
where $C$ is constant a corresponding to $C_{0}$, and the small parameter $\varepsilon$ characterizes the density of the finite-difference mesh, and $\theta$ ($|\theta|\to 1$) is a parameter. In case
$\theta=\frac{\sqrt{1+\varepsilon}}{2}(|C|\sqrt{\varepsilon}-\sqrt{\varepsilon C^{2}+4}).$ (2.8)
the scheme becomes exact, [^2] To simplify the calculations, in the next section we assume that $\theta=1$. As verified through previous calculations carried out by the authors, this assumption does not qualitatively affect the subsequent results, while it substantially simplifies the equations. i.e., its solutions at the grid nodes coincide with those of the corresponding ODE. As scheme (2.6), (2.7) is invariant, it admits the same algebra (1.5) as that of (1.4).
@@
$\begin{array}[]{c}\displaystyle J_{1}=\frac{C}{x_{+}-u}-\frac{\theta(u_{x}+1)}{\sqrt{u_{x}(x-u)(x_{+}-u_{+})}}=A,\\ \displaystyle J_{2}=\frac{Cx_{+}}{x_{+}-u}-\frac{\theta(x_{+}u_{x}+u)}{\sqrt{u_{x}(x-u)(x_{+}-u_{+})}}=B,\\ \displaystyle J_{3}=\frac{(x_{+}-x)(u_{+}-u)}{(x-u_{+})(x_{+}-u)}=\varepsilon.\end{array}$ (2.9)
@@
where $\rho$ is constant. Note that the constants $A$, $B$, and $C$ generally differ from the corresponding constants $A_{0}$, $B_{0}$, and $C_{0}$ in (1.6). While it should be possible to adjust them to match, this is not our objective here.
@@
$x_{n+1}=\frac{\varepsilon}{a(2\varepsilon+1)}\left\{\left((a^{2}+1)+\frac{a}{\varepsilon}+(a-1)\sqrt{a^{2}+\frac{2a(\varepsilon+1)}{\varepsilon}+1}\right)x_{n}\right.\\ \left.+\,b\left((a-1)+\sqrt{a^{2}+\frac{2a(\varepsilon+1)}{\varepsilon}+1}\right)\right\}.$
This is a linear equation solvable by standard methods. Substituting this and (2.14) into (2.6) taking into account (2.8) yields a quite cumbersome equation relating the values of $\varepsilon$, $a$, and $C$. This equation can be interpreted as a condition on the choice of the mesh density parameter $\varepsilon$.
@@ 3 Establishing a Finite-Difference Connection Through Bäcklund-type Transformations (Case $K=C^{2}=4$)
$\frac{(x-u_{-})(x_{-}-u)}{(u-u_{-})\sqrt{1+\varepsilon}}\left(\frac{x_{+}-u}{x_{-}-u}\frac{\sqrt{u_{x}}}{\sqrt{(x-u_{+})(x_{+}-u)}}-\frac{\sqrt{u_{\bar{x}}}}{\sqrt{(x-u_{-})(x_{-}-u)}}\right)=C.$ (3.1)
@@
$\frac{(u-x_{-})(u_{-}-x)}{(x-x_{-})\sqrt{1+\varepsilon}}\left(\frac{u_{+}-x}{u_{-}-x}\frac{\sqrt{x_{u}}}{\sqrt{(u-x_{+})(u_{+}-x)}}-\frac{\sqrt{x_{\bar{u}}}}{\sqrt{(u-x_{-})(u_{-}-x)}}\right)=\widetilde{C},$ (3.2)
@@
$x_{u}=\frac{x_{+}-x}{u_{+}-u},\qquad x_{\bar{u}}=\frac{x-x_{-}}{u-u_{-}},\qquad\widetilde{C}=\frac{\sqrt{\varepsilon}-\sqrt{1+\varepsilon}}{\sqrt{\varepsilon}+\sqrt{1+\varepsilon}}\,C.$